OpenStudy (anonymous):

Someone please help me? Write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. vertex at (–5, 0) and co-vertex at (0, 4)

OpenStudy (anonymous):

what do u kno how do so far

OpenStudy (anonymous):

Do you need help or nnot?

OpenStudy (anonymous):

i do need help...and no i dont know how to do it..

OpenStudy (mathmale):

While it's not strictly necessary, why not draw a picture of this ellipse? " a " is the length of the major axis, whereas " b " is the length of the minor axis. The standard equation of a horizontal ellipse centered at (0,0) is\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\]

OpenStudy (mathmale):

With this info., can you now determine the values of " a " and " b "? If so, substitute them into the equation above, and you'll have the desired equation of this ellipse.

OpenStudy (anonymous):

how do i find b?

OpenStudy (anonymous):

thanks for showing her/him @mathmale